Home‎ > ‎Reports‎ > ‎Crew100B‎ > ‎

Day01


Crew 100B Log Book for February 26, 2011


Photos of the day





Commander's Check-In Report

Balwant Rai Reporting


Crew Physical Status: Excellent, everyone is motivated for performing experiments

Time Departed/Returned from EVA:

0530 departure/ 0630 returned from EVA

Brief Narrative of Field Mission Results: Sample collection for bacterial activity; visited geology site; Saliva samples taken via Versi.SAL1 from Oasis diagnostics

EVA Data/Interpretations: As per reports of EVA

Engineering/Hab Maintenance: Oil, gas levels were checked. GreenHab was maintained.


Inventory: Food study inventory prepared for all crew members. Food pattern study for differences in vegetarian and non-vegetarian diet pattern starts today by Jasdeep Kaur

Support Requested: None

Miscellaneous: N/a


Commander's Report

Balwant Rai Reporting

Today we started with our first EVA and brief explanation of the medical experiments (including CASPER) that will be conducted during our stay.

We started the day with a briefing from crew 100A on all essential information for living at MDRS.

I started my medical experiment by taking saliva samples (via Versi Sal 1 from Oasis Diagnostics), vital parameters (via the Zephyr Bioharness), and analyzing heart rate variability (via NERV Express 4.2 software) from all crew members. I also performed a cognitive study (via CogState Research software) on all crew members.

The rest of the crew collected an inventory of all the food. All crew members also filled out the questionnaire related to the medical experiment and CASPER experiment.

The evening consisted of our first EVA, where Mattheiu, Jasdeep, and Quentin took out the ATV’s and collected rock/soil samples from a suspected habitat of extremophiles. They were guided by Rachel and Crystal, performing mission support. The direction in which the team went was at my suggestion. Today is a non-cooking day since we had a very busy day.

Tomorrow the plan is to do our second EVA (for geology and the medical experiments). We will also begin our daily medical experiments in the HAB. Radio telescope calibrations will also be performed tomorrow. The experiment or ergonomics will also begin tomorrow. Also, the food pattern study will begin tomorrow.


Engineering Report

Matthieu Ansart Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.75

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.44

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): Empty

Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 15

· Trailer Potable Water Tank Level (inches from bottom): 1/3 Full

· Hab Potable Water Tank Level (inches from bottom): 13

· Potable Water Meter Reading: 424946

Notes/Comments:


GreenHab:

· Condition of plants in Tank 1: Green water

· Condition of plants in Tank 2: Green things

· Condition of plants in Tank 3: Green things

· Greenhab notes/ comments:


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123973

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: No

· Fuel Consumed:

· Tire Status: Filled

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status:

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: Yes

· Oil Checked: No

· Fuel Consumed:

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: Yes

· Oil Checked: No

· Fuel Consumed:

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 70

· Actual temperature upstairs (Farenheit): 69

· Thermostat setting downstairs (Farenheit): 60

· Actual temperature downstairs (Farenheit): 58


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: All is well


General Engineering

1. Notes/Comments:

Clear


Science Report

Jasdeep Kaur Reporting

Time: 0700-Current

Crew Members:

Science Team: 3 crew members,

EVA mission supporters 1: 2 crew members

Site Location: Hab

Sample Collection Site: N 38.39203, W 110.79058

Transit Mode: ATV

Science Objectives: To locate the trace elements and bacterial activity.

Science Results: Soil sample is collected which has features of grayish white and cracked surface area, where most bacterial activity could be possible. It is preserved in a zip-lock bag for further analysis.

Further instructions: In ATV performance, scratches on helmet reduces visibility outside Hab, so need to be cleaned and defogged.

Green Hab Report: Visited the green hab and monitored all settings.


EVA 1 Report

Crystal Latham Reporting

EVA #1, 18:24-18:53

Crew Members: Matthieu Ansart, Quentin Bourges, Jasdeep Kaur

Site Location: (N 38.39203, W 110.79058)

Transit Mode: ATVs

Objectives: GeoMission 1: to collect rock/soil samples and measure radiation of rocks.

Results: Collected a sample of grayish-white rock and soil that was in an area that exhibited characteristics of water flowing, as well as ground fractures.

Problems: Members of the crew had trouble driving the ATVs for the first time. Also, mission support had problems hearing two of the three crew members, which was likely due to the fact that their microphones were too far away from their mouths. Crew members had problems seeing through helmets. Lessons Learned: Start missions prior to sunset. Make sure microphones do not move, so that crew members can still be heard by the mission support team. Have some practice riding the ATVs before participating in an EVA. Clean helmets thoroughly prior to EVAs so crew members can see well.


Chef's Report

Rachel Dompnier Reporting

1a) Today was a non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

No

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Stroganoff (with beef and Noodles) for some crew members and Santa Fe black beans and rice for other (part of a medical study of Jasdeep Kaur)

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

N/A

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

N/A

5) Lessons learned (if any)

N/A

6) Comments/questions for Kim and Jean, the food study investigators

N/A


Journalist Report

Quentin Bourges Reporting

Today was the arrival day at the MDRS. The crew 100A greeted us and gave us all the instructions we need for the next two weeks. After taking the traditional picture of everyone in front of the MDRS, crew 100A went back to Earth.

We had the opportunity to do the first EVA in which we took samples of rocks that will be analyzed. After that, the commander performed a medical experiment.

A briefing was also done in order to be sure of the role of each crew member.

Here is the beginning of a great experience!